OpenStudy (anonymous):

Be sure to distribute and collect like terms before you solve. 2(2x - 4) + 3x = -1

OpenStudy (anonymous):

2(2x - 4) + 3x = -1 4x - 8 + 3x = -1 7x - 7 = 0 7x = 7 x = 1.